OpenStudy (anonymous):

Which of the following is the first incorrect step? Step 4 Step 6 Step 7 Step 8

OpenStudy (anonymous):

Step 1 ax2 + bx + c = 0 Step 2 ax2 + bx = − c Step 3 x2 + b over a times x equals negative c over a Step 4 x2 + b over a times x plus b squared over 4 times a squared equals negative c over a plus b squared over 4 times a squared Step 5 x2 + b over a times x plus b squared over 4 times a squared equals negative 4 multiplied by a multiplied by c, all over 4 multiplied by a squared plus b squared over 4 times a squared Step 6 the square of the quantity x plus b over 2 times a equals b squared minus 4 multiplied by a multiplied by c, all over 4 times a squared Step 7 x plus b over 2 times a equals plus or minus square root of the quantity b squared minus 4 times a times c over 4 times a squared equals plus or minus square root of b squared minus 4 times a times c, all divided by 4 times a Step 8 x = −negative b over 2 times a plus or minus square root of b squared minus 4 times a times c, all divided by 2 times a Step 9 x = negative b plus or minus square root of b squared minus 4 times a times c, all divided by 2 times a

OpenStudy (anonymous):

Plz help someone FAN AND MEDAL

OpenStudy (anonymous):

its better if we put em to equation mode lol

OpenStudy (anonymous):

Yeah ik but teachers these days I swear they give them like this to me ._.

OpenStudy (anonymous):

lol irk XD

OpenStudy (anonymous):

step 4 looks like this \[\left[\begin{matrix}x^2+b \\ x+b^2 \\ 4a^2 \end{matrix}\right] = \left[\begin{matrix}-c \\ a+b^2 \\ 4a^2 \end{matrix}\right]\]

OpenStudy (anonymous):

idk how to 3 tier the fraction so i used matrix

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

Did you get the answer to this?

OpenStudy (anonymous):

that question successfully confused me

OpenStudy (anonymous):

back and yes it confused me too BLAME FLVS

OpenStudy (anonymous):

lol I did this question on a test yesterday

OpenStudy (anonymous):

Well then can you check your grade book and tell me what you got? XD

OpenStudy (anonymous):

Yeah hold on

OpenStudy (anonymous):

Step 4 is what I got

OpenStudy (anonymous):

XD ok thanks everyone for all

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

step 5 looks like \[\left[\begin{matrix}x^2+b \\ x+b^2 \\ 4a^2 \end{matrix}\right]=\left[\begin{matrix}-4ac \\ 4a^2+b^2 \\ 4a^2 \end{matrix}\right]\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i have no idea about step 6...dat english

OpenStudy (anonymous):

mostly just write em down and you can form that fraction

OpenStudy (anonymous):

lol and this is supposedly Algebra 1 ._.